OpenStudy (anonymous):

can some one help me finding the equivalent fraction on 1/6

OpenStudy (anonymous):

\[\frac{1}{6}=\frac{2}{12}=\frac{3}{18}=\frac{4}{24}=...\]

OpenStudy (anonymous):

thanks